OpenStudy (anonymous):

what is b equal to in 7b+4=9b-2

OpenStudy (anonymous):

7b - 9b = -2 - 4 -2b = -2 -2b/2 = -2/2 b = -2

OpenStudy (anonymous):

Wait oops I seem to have made a mistake. 7b - 9b = -2 - 4 - 2b = - 6 - 2b / -2 = -6 / -2 b = 3

OpenStudy (anonymous):

oh ok thank you very much

OpenStudy (anonymous):

No problem :)